CONFORMATIONAL INVESTIGATION OF A SERIES OF ISOLATED MODEL DIPEPTIDES

\underline{J. R. CABLE}, J. C. SHARP, and N. A. MILLER, Department of Chemistry and Center for Photochemical Sciences, Bowling Green State University, Bowling Green, OH 43403.

Model dipeptides based on single amino acids blocked at the amine with an acetyl group and at the carboxylic acid with an N-methylamide group have long served as tractable computational and experimental systems with which to model the relative stabilities of different peptide conformations. Replacing the N-methylamide blocking group with an N-phenylamide allows a very closely related set of compounds to be studied using electronic spectroscopy in supersonic jet expansions. Thus a series of compounds based on amino acids with small side chains, Ac-Gly-NHPh, Ac-Ala-NHPh, and Ac-Pro-NHPh, has been investigated. All evidence, including comparison to the vibronic spectra of the simpler N-phenylamides, formanilide and acetanilide, suggests that the dominant conformer in these model dipeptides adopts a C7 structure. Results from this study are compared to ab initio Hartree-Fock and MP2 calculations as well as to previous condensed phase spectroscopic studies on the N-methylamide-blocked systems.